OpenStudy (anonymous):

Please help medal and fan awarded Which line is represented by the function f(x) = 2x + 1? Line A Line B Line C Line D

OpenStudy (anonymous):

@Nnesha please help

Nnesha (nnesha):

where are lines

Nnesha (nnesha):

do you have any pic

OpenStudy (anonymous):

sorry I will get you a pic

OpenStudy (anonymous):

here you go thank you

Nnesha (nnesha):

ok this equation same like y = mx + b where m is slope which we use to rise and run and b is y-intercept which always draw on y-axis

OpenStudy (anonymous):

yes that is true but that is where Im confused what is the f(x)

Nnesha (nnesha):

f(x) same like y

OpenStudy (anonymous):

and the x is what

OpenStudy (anonymous):

Imean what is the x sorry

Nnesha (nnesha):

x is with slope is just the helper of m nothing have to do with it it helps us to see slope

OpenStudy (anonymous):

so the x is slope

Nnesha (nnesha):

no the number with x is slope

OpenStudy (anonymous):

what bout the x next to the 2 in the equation f(x)=2x+1

Nnesha (nnesha):

i was taking about that x

Nnesha (nnesha):

talking*******

Nnesha (nnesha):

were you asking about this one f(x) so its whole thing = y

OpenStudy (anonymous):

f(x)= y and the 2x is the slope

Nnesha (nnesha):

when question ask slope just say 2 x is not slope it is with slope

OpenStudy (anonymous):

oh ok so the answer would be B right?

OpenStudy (anonymous):

r u still there

OpenStudy (anonymous):

nvr mind

Nnesha (nnesha):

no b touches the y -axis at 0 but our y-intercept is one look your pic and tell which touche y-axis at 1

OpenStudy (anonymous):

the answer is A

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

thank you very much

Nnesha (nnesha):

no problem its my pleasure :)

OpenStudy (anonymous):

dont leave I might need more help and ill keep opening new questions so I caan give you more medals

OpenStudy (anonymous):

The first five terms of a sequence are shown below. 5, 9, 13, 17, 21 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 5n - 1; n ≥ 1 f(n) = 5n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1

OpenStudy (anonymous):

ignore the ????? in the diamonds

OpenStudy (anonymous):

@Flop can you please help

OpenStudy (anonymous):

k give me an attempt for the solution so I know where ur stuck

OpenStudy (anonymous):

I think the answer is B but I am not sure

Nnesha (nnesha):

sorry about that i wasn't know that you need more help

OpenStudy (anonymous):

oh its fine you dont have to

Nnesha (nnesha):

n represent term first plug 1 where n is then multiply by 5

Nnesha (nnesha):

you said b lets check yourself  is your 2nd term 11